, frictionless and weightless Piston was fitted into a cylinder containing gas this gas was allowed to expand from one litre to file against a constant pressure of 1 atmosphere in doing so 200 joule of heat was observed from the surrounding the change in the internal energy of the system is

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Work done in expansion=\int PdV=P\Delta Vat constant pressure.

=\implies W=10^5 (V_f-10^-3)joules

From conservation of energy,Q=\Delta U+W
\implies \Delta U=Q-W=(+200)-(10^5(V_f-10^{-3}))

The final value of the volume is not given here. Put in to obtain the answer.

Regards.